OpenStudy (anonymous):

If sinhx=8/15, find the values of the other hyperbolic functions at x. Anyone know the formula to this?

Directrix (directrix):

(coshx)^2 - (sinhx)^2 = 1 (coshx)^2 - (8/15)^2 = 1 (coshx)^2 = 289/225 (coshx) = +/- 17/15 See other equations at the link below: http://www.sosmath.com/trig/hyper/hyper01/hyper01.html